OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth. Show all work.

OpenStudy (anonymous):

3% per year?

OpenStudy (anonymous):

if so, then 2500*1.03^n=5000 => 1.03^n = 2 => n*log1.03=log2 => n= log2/log1.03 = 23.4 months.

OpenStudy (anonymous):

previously was for per month... if per year: 23.4 yr * 12 = 281 months.

OpenStudy (anonymous):

Thank you.